If vector x * vector b = vector c * vector b and vector x . vector c =0 then find vector x
1
Expert's answer
2020-05-25T21:23:58-0400

"\bar{x}\times\bar{b}=\bar{c}\times\bar{b}=>\begin{matrix}
 or\ \bar{x}=\bar{c} \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
 or\ \bar {b}=0\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
 or\ \bar{x}=\bar{c}=0 \ and\ \bar{b}\not=0
\end{matrix}"

Given "\bar{x}\cdot\bar{c}=0"

If "\bar{b}\not=0," then "\bar{x}=0"